Circular No. 7671

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001de IN UGC 12089 T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics, report that a spectrum of SN 2001de (cf. IAUC 7664), obtained by P. Berlind on July 24.42 UT with the Mt. Hopkins 1.5-m telescope (+ FAST spectrograph), shows it to be a type-Ia supernova, ∼ 1 week after maximum light. Adopting the NED recession velocity of 9313 km/s for the host galaxy, the supernova expansion velocity is ∼ 10 600 km/s for Si II (rest 635.5 nm). The spectral-feature age of the supernova (Riess et al. 1997, A.J. 114, 722) is 6 ± 2 days after maximum light.

N. V. Borisov, A. N. Burenkov, V. V. Vlasuk, and A. G. Pramsky, Special Astrophysical Observatory, report on their observations of SN 2001de obtained with the 6-m telescope (+ long-slit spectrograph + CCD camera) on July 22.04 UT: “A full reduction of the CCD spectrogram (range 360–810 nm, dispersion 0.24 nm/pixel) shows that this object is indeed a supernova. The spectral features are consistent with a peculiar type-Ia object. The usual Si II feature (635.5 nm) is much less deep than in normal type-Ia supernovae, if compared with all of the other spectral lines. The spectrum is dominated by P-Cyg profiles, superimposed on a slightly blue continuum. The most prominent features are identified as Ca II H and K, Si II, Mg II, and S II. The expansion velocities measured from the minima of Ca II H and K, Mg II 448.1-nm, and Si II 635.5-nm are 13 800, 11 400, and 11 000 km/s, respectively. The recession velocity of the parent galaxy, deduced from the emission peak of the Si II 635.5-nm line, is 9700 km/s.”

V1333 AQUILAE C. Bailyn, S. Tourtellotte, and L. Walkowicz, Yale University; and

D. Gonzalez and J. Espinoza, Cerro Tololo Interamerican Observatory (CTIO), report that the current outburst of Aql X-1 reported on IAUC 7646 continues to show very unusual characteristics: “We have continued daily R-band observations of Aql X-1 (weather permitting) with the Yale 1-m telescope (+ ANDICAM) at CTIO. We find that, for the past several weeks, the source has varied from 0.8 to 1.3 mag brighter than quiescence. Since our measurements in both quiescence and outburst include light from the near neighbor, V1333 Aql itself must be 2–3 mag brighter than quiescence. During this time, the RXTE All Sky Monitor shows x-ray flux levels of < 3 counts/s. Thus this outburst is quite different from normal outbursts of this source, which are accompanied by x-ray emission an order of magnitude greater than has been observed during this outburst, and from previously reported ‘mini-outbursts’, which last only a few days.”

2001 July 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green